Which equation represents the line that passes through the points (-4,-1) and (-1,-7)?
Answer by rfer(16322)

m=(-7--1)/(-1--4)
m=-6/3
m=-2
---------------
y=mx+b
-1=-2*-4+b
-1=8+b
b=-9
-------------
y=-2x-9